OpenStudy (devonhoward15):

To solve the equation below by completing the square, what is your first step? 2x^2+12x=32

OpenStudy (devonhoward15):

@mathstudent55 @calculusxy

OpenStudy (devonhoward15):

@ganeshie8

OpenStudy (paigejboo):

^2

OpenStudy (devonhoward15):

@mathstudent55

OpenStudy (peachpi):

I'd factor the 2

OpenStudy (paigejboo):

I gave you the answer?

OpenStudy (devonhoward15):

i dont have either of those as a anwser

OpenStudy (paigejboo):

Do the first one!

OpenStudy (devonhoward15):

@Ms-Brains

OpenStudy (peachpi):

multiplying by 1/2 is the same thing as dividing by 2

OpenStudy (devonhoward15):

@563blackghost

563blackghost (563blackghost):

Let me see. Hold on for sec.

OpenStudy (devonhoward15):

ok

OpenStudy (mathstudent55):

You need x^2, not 2x^2. To get that, you need to divide the entire equation by 2. Multiplying by 1/2 is the same as dividing by 2.

OpenStudy (devonhoward15):

ok thanks

563blackghost (563blackghost):

So we would do.... \(\Huge{\frac{2x^{2} + 12x~ =~ 32}{2}}\)

563blackghost (563blackghost):

So it would simplify to \(\Huge{x^{2} + 6x = 16}\) Which we want. Now we would subtract 16 from the right side and add -16 to the left side.... \(\Huge{x^{2} + 6x -16=0}\) This is what we want. Now we determine what x could be.

OpenStudy (mathstudent55):

After you have the x^2 term as simply x^2 with no coefficient, keep the x^2 term and the x term on the left side, and move the number to the right side.

OpenStudy (mathstudent55):

\(\Huge{x^{2} + 6x = 16}\) Now to complete the square, we take half of the x-term coefficient, and we square it. Half of 6 is 3. Square 3 to get 9. We add this to both sides: \(\Huge x^{2} + 6x \color{red}{+9}= 16 \color{red}{+9}\) That was the step that completed the square. Now we can simplify the right side, and rewrite the left side as the square of a binomial.

OpenStudy (mathstudent55):

\(\Huge (x + 3)^2 = 25\)